OpenStudy (anonymous):

choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3.

OpenStudy (anonymous):

y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

use the formula (y-y1)=m(x-x1) where m is the slope and x1,y1 is the given point!!

OpenStudy (anonymous):

so therefore your answer is y-6=-3(x+1)!

OpenStudy (anonymous):

thank you !! can u help me with one more ?

OpenStudy (anonymous):

yeah sure!!

OpenStudy (anonymous):

okay here u have to use the formula \[\frac{ y-y _{1} }{ y _{2}-y _{1} }=\frac{ x-x _{1} }{ x _{2}-x _{1} }\] where x1,y1 is one point and x2,y2 is the 2nd point!!